OpenStudy (anonymous):

w+4=-10 The answer that i got is w=6 Is that right

OpenStudy (anonymous):

w=-14

OpenStudy (anonymous):

What about this one n-8=5

OpenStudy (anonymous):

w+4=-10 subtracting 4 on both sides w+4-4=-10-4 w=-14

OpenStudy (anonymous):

What about this one n-8=5

OpenStudy (anonymous):

n-8+8=5+8 adding 8 on both sides n=13

OpenStudy (anonymous):

r-4=-18 ?

OpenStudy (anonymous):

Emra add 4 on both sides of the equation. Try to do it yourself. and let me know.

OpenStudy (anonymous):

Y-3=-1 This is what i got is it right y=4

OpenStudy (anonymous):

No, Emra you are grouchy.........:)

OpenStudy (anonymous):

Y-3=-1 This is what i got is it right y=4

OpenStudy (anonymous):

r-4=-18 r-4+4=-18+4 r=-14

OpenStudy (anonymous):

y-3=-1 what did you get ^ ^ Thats wrong